# Hyperband-based Bayesian Optimization for Black-box Prompt Selection

### v1

@@ 5.2 Analysis of the performance for each LLMs
As demonstrated in Section 5.1, TRIPLE-SH emerges as the strongest competitor. To assess whether HbBoPs’s improvements are consistent across different LLMs, we present in Table 2 the median relative improvement over the ten benchmark tasks for each LLM, along with the interquartile range in parentheses.
The table reveals that HbBoPs consistently outperforms TRIPLE-SH in terms of both anytime validation and test performance. For instance, when using Claude 3 Haiku, the average test error is reduced by a median factor of $0.066$ and $0.027$ at $0.25$ and $0.50$ of the total budget, respectively.
While we observe positive improvements over TRIPLE-SH on the validation set with a full budget, these gains are less pronounced on the test set. This discrepancy might be attributed to both methods successfully identifying the optimal prompt for each benchmark when having evaluated more prompts on more validation instances but therefore also having made more LLM calls.
@@ 5.3 Ablation study
Figure 2: Anytime normalized error of the best prompt found by each ablation variant and RS as a baseline, averaged over different benchmark scenarios. Lower is better. Log scale. Ribbons represent standard errors.
Table 3: Normalized validation and test error using HbBoPs with distinct encoder models for different fractions of total LLM calls averaged over all 30 benchmarks. Standard errors in parentheses.
@@
To better understand the contributions of the individual components in HbBoPs, we conduct a comprehensive ablation study. Figure 2 reports the average anytime normalized validation and test errors of the best prompt found by systematically removing specific components of our HbBoPs such that we can quantify their importance.
Our analysis focuses on four key aspects: the use of a GP with a deep kernel, the incorporation of a structural-aware deep kernel, the integration of Hyperband for multi-fidelity scheduling, and the final proposed method. Results demonstrate the incremental improvements achieved by each component. We focus on the validation error as shown in Figure 2(a) to describe results as improvement here is a direct consequence of the change of components, but, we report the test error in Figure 2(b).
We first observe that using a GP with a deep kernel (BoPs (non structural-aware DK-GP)) in a full-fidelity BO setting on average improves over vanilla BO by $11\%$ and $38\%$ at $0.5$ and $1.0$ budget with respect to the average normalized validation error. These improvements highlight the importance of properly handling the high-dimensional embedded space. Building upon this, the structural-aware deep kernel (BoPs (structural-aware DK-GP)) further enhances performance by $9\%$ and $13\%$ at $0.5$ and $1.0$ budget, respectively, demonstrating the value of directly incorporating structural information into the deep kernel.
The integration of Hyperband for multi-fidelity scheduling (HbPs using a random proposal mechanism) provides orthogonal boosts to both anytime and final performance. We observe improvements of the average normalized validation error of $47\%$ and $11\%$ at $0.5$ and $1.0$ budget, respectively, over the full-fidelity BoPs (structural-aware DK-GP). Finally, our complete HbBoPs further increases performance, achieving a $21\%$ improvement over HbPs at $0.5$ budget and $31\%$ at $1.0$ budget. Notably, when compared to our starting point of vanilla BO, HbBoPs demonstrates a substantial $66\%$ improvement at $0.5$ budget and $67\%$ at $1.0$ budget.
@@ 5.4 Analysis of varying the encoder
As HbBoPs relies on embeddings of prompts, we conduct a sensitivity analysis to evaluate the impact of different encoder models on the performance of HbBoPs. While our primary results were obtained using BERT’s (Devlin et al., 2019) [CLS] token embedding, we extend our analysis to include two more encoder models that are MPNet (Song et al., 2020) and DistillRoBERTa (Liu et al., 2019). For each encoder model, we rerun all benchmarks with HbBoPs.
We report in Table 3 the average normalized validation and test error for each encoder when used within HbBoPs for different fractions of total LLM calls over all 30 benchmark scenarios. Results demonstrate that the average normalized validation and test performance of HbBoPs remains within a similar range across all encoder models, which suggests that HbBoPs is robust to the choice of encoder. This robustness can be expected, as none of the encoder models were specifically trained or fine-tuned for our task of predicting prompt performance. The effectiveness of HbBoPs stems from its ability to learn a mapping from prompts to performance through the DK-GP, provided that the raw embedding features can semantically differentiate between prompts and carry relevant information.
@@ 6 Related Work
6.1 Automatizing Prompt Engineering
@@
As the construction and evaluation of a large number of prompts on the validation set can be a huge bottleneck, recent work has been concerned with the general topic of automatizing prompt engineering. This work can be classified into prompt optimization, i.e., automating the creation of prompts (Prasad et al., 2022; Sun et al., 2022a; Zhou et al., 2022; Xu et al., 2022; Diao et al., 2022; Chen et al., 2023; Lin et al., 2023; Fernando et al., 2023; Pryzant et al., 2023; Guo et al., 2023; Pan et al., 2023; Schnabel & Neville, 2024; Shen et al., 2023; Hu et al., 2024), and prompt selection, i.e., finding the best prompt within a finite set (Wu et al., 2024; Shi et al., 2024; Opsahl-Ong et al., 2024; Do et al., 2024).
@@
Another dimension to categorize the related literature is given by the white-box vs. black-box setting. The white-box setting assumes access to the LLM, so that gradient-based methods for prompt optimization or selection are applicable (Shin et al., 2020). The black-box setting assumes no access to the LLM which only returns an output given input (Sun et al., 2022b; Diao et al., 2022).
@@
Finally, another differentiation is given by the static and dynamic setting. The goal of the static setting (Wu et al., 2024; Shi et al., 2024; Khattab et al., 2023) is to obtain a single prompt that in expectation performs well for all instances during test time. In contrast, the goal of the dynamic setting (Zhang et al., 2022; Do et al., 2024; Xu et al., 2024) is to select a prompt for each test instance (Rubin et al., 2021).
6.2 Static Black-box Prompt Selection
@@
MIPROv2is DSPy’s (Khattab et al., 2023) state-of-the-art teleprompter for joint instruction and few-shot exemplar selection. MIPROv2 searches over a finite set of candidate prompts by combining instructions with few-shot exemplars (which DSPy first constructs automatically). While MIPROv2 is a variant of BO, it learns a surrogate Tree-Structured Parzen Estimator (TPE; Bergstra et al. 2011) based on the categorical indices of instructions and exemplars ($\text{ID}_{i}$ and $\text{ID}_{e}$) that compose a prompt. A downside is that learning a surrogate model based on indices does not use any semantic information of the prompts which may result in a sub-optimal mapping from prompts to their performance. Moreover, MIPROv2 does not directly address query efficiency. Indeed, while DSPy can be configured to use a smaller random subset of the validation data to evaluate prompts, this risks sub-optimal selection due to noisy performance estimates (see also Appendix B.
@@
EASE proposed by Wu et al. (2024) mainly focuses on few-shot exemplar selection. It uses NeuralUCB (NUCB; Zhou et al. 2020) with prompt embeddings as features, allowing for sequential evaluation of promising prompts based on the UCB criteria. EASE’s main contribution, however, is to make the combinatorial problem of selecting examples to build the few-shot exemplar from a larger training set computationally feasible. It prunes the candidate space using an optimal transport inspired heuristic before applying UCB. However, EASE is affected by query inefficiency since it evaluates prompts on all validation instances (or a random subset again risking sub-optimal selection).
TRIPLE is a class of algorithms for black-box prompt selection recently proposed by Shi et al. (2024) to address query-efficiency. TRIPLE solves the problem of black-box prompt selection via a multi-armed bandit approach. In details, the authors rely on existing bandit algorithms such as Successive Halving (SH; Karnin et al. 2013 or Generalized Successive Elimination (GSE; Azizi et al. 2021). These algorithms accelerate prompt evaluation by discarding poorly performing prompts early, thereby avoiding the evaluation of all prompts on all validation instances during the selection process. TRIPLE-SH and TRIPLE-GSE have the following limitations: TRIPLE-SH is sensitive to the initial evaluation budget, potentially discarding promising prompts based on noisy performance estimates. This is especially problematic when evaluating many prompts with a limited budget which often is the case when jointly searching over instructions and few-shot exemplars. TRIPLE-GSE attempts to mitigate this by modeling the expected prompt performance based on embeddings projected to a lower dimensional space via a random projection and point-wise performance observations of the prompts. However, GSE has not been analyzed for structured bandit problems beyond the generalized linear setting (Azizi et al., 2021).
@@ 7 Conclusion
Existing methods applicable to static joint instruction and few-shot exemplar selection for black-box LLMs suffer from either not being sample-efficient, i.e., they evaluate all prompts on at least some validation instances, or not being query-efficient, i.e., they use all validation instances during the evaluation of each prompt. In this work we have proposed HbBoPs, a both sample- and query-efficient method tailored to joint instruction and few-shot exemplar selection. HbBoPs relies on a structural-aware deep kernel Gaussian Process that models the downstream performance of prompts which is used to propose promising, yet not evaluated prompts during the selection process. In addition, HbBoPs employs a Hyperband scheduler that governs the number of validation instances prompts are evaluated on. Our extensive experiments have demonstrated that HbBoPs improves upon existing baselines and state-of-the-art competitors in the limited budget regime while showing strong anytime performance. Future work might expand HbBoPs to a multi-objective setting. As for example the inclusion of many examples in a prompt induces longer response latency, the trade-off between the number but also concrete selection of examples for a prompt give rise to a multi-objective optimization problem of performance and response latency. Ultimately, we believe that our HbBoPs presents a significant step forward in making prompt selection for black-box LLMs more accessible and efficient, accelerating the usage of these models across a broader range of applications.
@@
Appendix A Illustrations of the structural-aware Deep Kernel GP
As mentioned in Section 3.1, unsupervised dimensionality reduction techniques such as PCA or random projections will not result in a lower latent dimension that is aligned with the downstream performance of prompts. To illustrate this, we perform the following experiment. We collect the validation performance of 250 prompts on the GSM8K task using LLAMA3 8B Instruct as LLM. We embed each prompt using the [CLS] token embedding of BERT ($768d$). We then split the prompts $p_{1},\ldots,p_{250}$ and their corresponding validation errors $v_{1},\ldots,v_{250}$ in a train (80%) and test set (20%). Using the train set we perform a PCA retaining ten principle components as features. Moreover, we train our structural-aware deep kernel GP introduced in Section 3.2 on the training set and extract the $10d$ latent features from the output of the feature extractor $\phi_{\left(\phi_{enc(i)},\phi_{enc(e)}\right)}$. We visualize the raw $768d$ embedding features of prompts, the $10d$ PCA features and the $10d$ deep kernel features for the training set using a two component t-SNE (Van der Maaten & Hinton, 2008) in the top row of Figure 3. The $x$- and $y$-axis represent the two t-SNE components, whereas color indicates the validation error of prompts (lighter being better). We can see that for both the raw embedding features and the PCA features, it is difficult to visually make out any meaningful clusters or structure how closeness in feature space relates to closeness in performance space. For the deep kernel features, however, we can see that the feature space is nicely aligned with the performance space (well performing prompts being closer together with a continuous transition into poorer performing prompts) - of course, this is on the training split on which the GP has been trained on so it is not too surprising. However, when looking at the test split (that was neither used to perform the PCA, nor to train the GP) in the bottom row of Figure 3, we can see that similar conclusions as for the train split hold. The latent representation the feature extractor of the deep kernel has obtained during training does generalize and it has effectively learned its own low-dimensional embedding that is in relevance to the downstream task.
Figure 3: Visualization using a two component t-SNE of the $768d$ BERT [CLS] token embeddings of prompts. Left: Raw unprocessed features. Middle: Features of a ten component PCA solution. Right: Latent features ($10d$) from the feature extractor of our structural-aware deep kernel GP. Top row: Train split. Bottom row: Test split. Color indicates the performance of prompts for LLAMA3 8B Instruct on GSM8K.
@@ Appendix B On the Generalization from Validation to Test
When performing black-box prompt selection, we evaluate prompts on a validation set and iteratively improve over the current best prompt by trying to identify a better one. While progress on the validation set is expected (i.e., if we perform full-fidelity evaluations using the same validation instances for each prompt, the validation error of the incumbent will be monotonically decreasing as optimization progresses), it must not necessarily be the case that we also improve performance on a separate test set of instances, i.e., the prompt identified as being validation optimal might not be optimal on the test set.
In this section we provide some additional results regarding generalization from validation to test. Recall that many methods for black-box prompt selection (e.g., EASE and MIPROv2) are by design not query-efficient but evaluate all prompts on all validation instances or a random subset (e.g,. Wu et al. 2024 used validation sets containing as few as 20 validation instances in their benchmark experiments). We will now empirically demonstrate that using small random subsets during optimization is not a sensible choice, because this will result in increased variance of the estimate of the validation error which prevents us from making correct decisions on the validation set and can result in generalization issues when moving from the validation set to the test set.
To illustrate this, we perform the following experiment. We collect the validation and test error (according to the splits described in Table 5) of 250 prompts on the GSM8K task using LLAMA3 8B Instruct as the LLM. We now vary the number of validation instances used to evaluated the performance of a prompt via bootstrapping, using $k=10,50,100,500$ instead of the original $n_{\mathrm{valid}}=1319$ validation instances. Note that for the test performance we always use the full test set. In Figure 4 we provide scatter plots of the validation and test errors of the prompts with mean validation errors obtained via bootstrapping using $k=10,50,100,500$ validation instances vs. validation errors obtained on the full validation set of $1319$ instances. We perform $1000$ bootstrap replicates.
@@
We can observe that if we use too few validation instances ($k=10$ but also $k=50$ and $k=100$ to some extent) even if we select the validation optimal prompt, its test performance can be far from optimal, because noise dominates the estimate of the validation error.
In Figure 5 we provide box-plots of the bootstrapped variance estimates of the mean validation error over the prompts when using smaller validation sets.
Figure 5: Box-Plots of the bootstrapped variance estimates of the mean validation error of 250 prompts evaluated with LLAMA3 8B Instruct on GSM8K varying the number of validation instances used to estimate the mean validation error.
As expected, the variance of the mean validation error can be substantial when using few validation instances. Note that bootstrap results are generally in line with what one would expect when assuming the point-wise (binary) loss (based on exact match as a scoring function) of a prompt on a validation instance being Bernoulli distributed with success probability $p$ of a loss of 00 as it then follows that the validation error is Binomial distributed based on the number of validation instances $n_{\mathrm{valid}}$ and its average validation error as a plug-in estimate for the success probability $p$, and therefore we obtain a variance of the estimated validation error of $\nicefrac{{p(1-p)}}{{n_{\mathrm{valid}}}}$; e.g., if a prompt has a true success probability of a loss of 00 of $p=0.5$ we expect a variance of the estimated validation error of $\nicefrac{{0.5^{2}}}{{10}}=0.025$ when using $10$ validation instances to estimate its validation error.
This has serious practical implications depending on the variation of the performance of prompts on a downstream task. Because if the true performance of many prompts is similar, we cannot hope to be able to tell them apart based on their estimated validation errors as noise dominates the signal when using too few validation instances. Also, for benchmarking methods for black-box prompt selection this is highly relevant as when using too few validation instances, we cannot know if a generalization gap from the validation set to the test set is solely a result of noisy performance estimates or also due to internal mechanisms of a method (such as the optimal transport inspired heuristic employed by EASE to only consider examples that are similar to the validation set) which may further result in overfitting to the validation set.
The results we have presented here further provide justification for using a multi-fidelity scheduler over the validation instances during prompt selection, such as HB. Poor performing prompts can be differentiated using few validation instances, however, well-performing prompts need to be evaluated on larger sets so that one can effectively tell their validation performances apart.

### v2

@@ 5.2 Analysis of the Performance for each LLM
As shown in Section 5.1, TRIPLE-SH emerges as the strongest competitor. To assess whether HbBoPs’s improvements over TRIPLE-SH are consistent across different LLMs, we present in Table 2 the median relative improvement over the ten benchmark tasks for each LLM.
The table reveals that HbBoPs consistently outperforms TRIPLE-SH in terms of both anytime validation and test error. For instance, when using Claude 3 Haiku, the average test error is reduced by a median factor of $0.066$ and $0.027$ at $0.25$ and $0.50$ of the total budget, respectively. While we further observe positive improvements over TRIPLE-SH on the validation set with a full budget, these gains are less pronounced on the test set. One possible reason for this discrepancy is that both methods, when given a sufficient budget, successfully identify prompts that have low or optimal validation error. However, optimal validation error does not necessarily guarantee optimal test error. We provide further analysis of the generalization gap between validation and test performance that is influenced by the size of the validation set in Appendix B.
@@ 5.3 Ablation Study
Figure 2: Normalized error (log scale) of the best prompt per HbBoPs ablation variant, RS, and vanilla BO, averaged over benchmarks. Lower is better. Ribbons represent SE.
Table 3: Normalized validation and test error of HbBoPs with different encoders at different fractions of total LLM calls averaged over all 30 benchmarks. SE in parentheses.
@@
To better understand the contributions of the individual components of HbBoPs, we conduct a comprehensive ablation study. Our ablation focuses on four key aspects: the use of a GP with a deep kernel (DK-GP), the incorporation of a structural-aware DK-GP, the integration of HB for multi-fidelity scheduling, and the final HbBoPs. We aim to answer the following main research questions: (RQ1) Does a structural-aware DK-GP improve over a non-structural-aware DK-GP and a vanilla GP? (RQ2) Does multi-fidelity scheduling with HB improve over full-fidelity methods? (RQ3) Does combining our structural-aware DK-GP with HB improve over HB with a random proposal?
Figure 2 presents the average anytime normalized validation and test errors of the best prompt found by systematically removing specific components of our HbBoPs such that we can quantify their importance and answer above research questions. We focus on the validation error as shown in Figure 2(a) to describe results, as improvement of the validation error is a direct consequence of the change of components.
First, we observe that using a DK-GP on prompts embedded as a block of text (BoPs (non structural-aware DK-GP)) in a full-fidelity setting improves over vanilla BO by $11\%$ and $38\%$ at $0.5$ and $1.0$ budget with respect to the average normalized validation error. This highlights the importance of handling the high-dimensional embedded space properly. The structural-aware deep kernel (BoPs (structural-aware DK-GP)) further enhances performance by $9\%$ and $13\%$ at $0.5$ and $1.0$ budget, demonstrating the value of directly incorporating structural information into the GP which answers (RQ1). Note that the structural-aware DK-GP improves over HDBO by $19\%$ and $8\%$ with respect to final average normalized validation and test error.
The integration of HB for multi-fidelity scheduling (HB using a random proposal mechanism) provides orthogonal boosts to both anytime and final performance. We observe improvements of $47\%$ and $11\%$ at $0.5$ and $1.0$ budget over the full-fidelity BoPs (structural-aware DK-GP) answering (RQ2).
Our complete HbBoPs further increases performance, achieving a $21\%$ improvement over HB at $0.5$ budget and $31\%$ at $1.0$ budget, answering (RQ3). Compared to our starting point of vanilla BO, HbBoPs demonstrates a substantial $66\%$ improvement at $0.5$ budget and $67\%$ at $1.0$ budget. For additional statistical analyses, we refer to Appendix E.
@@ 5.4 Analysis of Varying the Encoder
As HbBoPs relies on embeddings of prompts, we conduct a sensitivity analysis to evaluate the effect of different encoder models on the performance of HbBoPs. While our primary results were obtained using BERT’s (Devlin et al., 2019) [CLS] token embedding, we extend our analysis to include two more encoder models that are MPNet (Song et al., 2020) and DistillRoBERTa (Liu et al., 2019). For each encoder model, we rerun the entire set of benchmarks.
We report in Table 3 the average normalized validation and test error for each encoder when used within HbBoPs for different fractions of total LLM calls over all 30 benchmark scenarios. Results show that HbBoPs maintains consistent validation and test error across all encoder models, indicating robustness to the choice of encoder. This is expected, as none were specifically fine-tuned for predicting prompt performance. HbBoPs’s effectiveness stems from its ability to learn a mapping from prompts to performance through the structural-aware DK-GP, provided the embeddings capture meaningful distinctions between prompts.
@@ 6 Related Work
Automating Prompt Engineering.
@@
Recent work has been concerned with the general topic of automating prompt engineering. This work can be classified into prompt optimization, i.e., automating the creation of prompts (Prasad et al., 2023; Sun et al., 2022a; Zhou et al., 2023; Xu et al., 2022; Diao et al., 2023; Chen et al., 2024; Lin et al., 2024; Fernando et al., 2024; Pryzant et al., 2023; Guo et al., 2024; Pan et al., 2024; Schnabel & Neville, 2024; Shen et al., 2023; Hu et al., 2024), and prompt selection, i.e., finding the best prompt within a finite candidate set (Wu et al., 2024; Shi et al., 2024; Opsahl-Ong et al., 2024; Do et al., 2024).
@@
Another dimension to categorize the related literature is given by the white-box vs. black-box setting. The white-box setting assumes access to the LLM, so that gradient-based methods for prompt optimization or selection are applicable (Shin et al., 2020). The black-box setting assumes no access to the LLM which only returns output given input (Sun et al., 2022b; Diao et al., 2023).
@@
Finally, another differentiation is given by the static vs. dynamic setting. The goal of the static setting (Wu et al., 2024; Shi et al., 2024; Khattab et al., 2024) is to obtain a single prompt offline that in expectation performs well for all instances during test time. In contrast, the goal of the dynamic setting (Zhang et al., 2023; Do et al., 2024; Luo et al., 2024) is to select a prompt for each test instance (Rubin et al., 2022) in an online fashion.
Static Black-box Prompt Selection.
Our work falls into the category of static black-box prompt selection. We summarize existing works below.
@@
MIPROv2 (Opsahl-Ong et al., 2024) is DSPy’s (Khattab et al., 2024) state-of-the-art teleprompter for joint instruction and few-shot exemplar selection. It searches over a finite set of candidate prompts by combining instructions with few-shot exemplars (which DSPy first constructs automatically). The method is a variant of BO using a Tree-structured Parzen Estimator (TPE; Bergstra et al. 2011) based on the categorical indices of instructions and exemplars ($\text{ID}_{i}$ and $\text{ID}_{e}$) that compose a prompt. A downside is that learning a surrogate model based on indices does not use any semantic information of prompts, which may result in suboptimal predictive performance. Moreover, MIPROv2 does not directly address query-efficiency. While DSPy can be configured to use a smaller random subset of the validation set to evaluate prompts, this risks suboptimal selection due to noisy performance estimates (see also Appendix B).
@@
EASE proposed by Wu et al. (2024) mainly focuses on few-shot exemplar selection. It uses NeuralUCB (NUCB; Zhou et al. 2020) with embeddings of prompts as blocks of text as features, allowing for sequential evaluation of promising prompts based on the UCB criteria. EASE’s main contribution is to make the combinatorial problem of selecting examples to build the few-shot exemplar from an extensive training set computationally feasible. It prunes the candidate space using an optimal transport inspired heuristic before applying UCB. EASE is affected by query-inefficiency since it evaluates prompts on all validation instances (or a random subset, again risking suboptimal selection).
TRIPLE proposed by Shi et al. (2024) is a class of query-efficient algorithms for static black-box prompt selection using a multi-armed bandit approach. It makes use of Successive Halving (Karnin et al., 2013) or Generalized Successive Elimination (Azizi et al., 2022) to accelerate prompt evaluation by discarding poor-performing prompts early, reducing the need to evaluate all prompts on all validation instances. However, TRIPLE-SH is sensitive to the initial evaluation budget and may prematurely discard promising prompts due to noisy performance estimates. TRIPLE-GSE tries to mitigate this by (non-linear) modeling of expected prompt performance using embeddings of prompts as blocks of text projected to a lower-dimensional space. While this approach introduces flexibility, the Generalized Successive Elimination algorithm has been formally analyzed only in the generalized linear setting (Azizi et al., 2022). Moreover, both TRIPLE-SH and TRIPLE-GSE begin by evaluating all prompts, whereas our HbBoPs employs a sample-efficient BO proposal to select candidate prompts for evaluation.
@@ 7 Conclusion
We introduced HbBoPs, a method for static black-box prompt selection in which prompts are composed of instructions and few-shot exemplars. HbBoPs employs a structural-aware deep kernel Gaussian Process to model the downstream performance of prompts based on separate embeddings of instructions and exemplars. This enables the identification of promising, unevaluated prompts during the selection process, making HbBoPs highly sample-efficient. Furthermore, HbBoPs integrates Hyperband as a multi-fidelity scheduler that governs the number of validation instances used for prompt evaluation, ensuring query-efficiency. In extensive experiments, we have demonstrated that HbBoPs improves upon baselines and state-of-the-art competitors in the limited budget regime while showing strong performance at any stage of the selection process.
While HbBoPs demonstrates strong performance, some limitations remain. Our method depends on embeddings from pre-trained encoders. Although we have demonstrated that HbBoPs is largely robust to the concrete choice of encoder model, obtaining embeddings induces some minor computational overhead. Additionally, our analysis focused on prompts composed of instructions and few-shot exemplars. While these components are highly relevant in practice, prompts may also include additional elements such as output guidance, formatting constraints, or other structural cues, which remain unexplored in this work. Nevertheless, the use of a deep kernel Gaussian Process as a surrogate model makes our framework, in principle, flexible enough to incorporate such additional prompt parameters by, for example, including categorical variables for formatting styles (e.g., JSON, Markdown).
In our experiments, we evaluated HbBoPs in a static setting with a fixed set of candidate prompts generated a priori. This enabled fair comparisons across baselines and methods operating in the static black-box prompt selection setting. However, we emphasize that HbBoPs can also be applied in more flexible prompt optimization settings where the candidate pool evolves over time. For example, it could be integrated with mutation-based prompt generation strategies (Fernando et al., 2024), where new prompts are generated iteratively, or with similar end-to-end prompt optimization pipelines (Pryzant et al., 2023; Yang et al., 2024).
Future work could extend HbBoPs to multiple objectives. Selecting more examples to include in a prompt may enhance performance but also increases response latency. Balancing the number and composition of examples in a few-shot exemplar introduces a trade-off between performance and efficiency, giving rise to a multi-objective optimization problem.
@@
Appendix A On the Latent Space of the Structural-Aware Deep Kernel Gaussian Process
As described in Section 3.1, unsupervised dimensionality reduction techniques such as PCA or random projections will not result in a lower-dimensional latent representation of prompt embeddings that is aligned with the downstream performance of prompts. To illustrate this, we perform the following experiment: We collect the validation error of 250 prompts on GSM8K (according to the splits described in Table 5) using LLAMA3 8B Instruct as LLM. We embed each prompt using the [CLS] token embedding of BERT ($d=768$). We then split the prompts $p_{1},\ldots,p_{250}$ and their corresponding validation errors $v_{1},\ldots,v_{250}$ in a train (80%) and test set (20%). Using the train split, we perform a PCA and retain $10$ principal components as features. Moreover, we train our structural-aware DK-GP introduced in Section 3.2 on the training split and extract the $10$ latent features from the output of the feature extractor $\phi_{\left(\phi_{enc(i)},\phi_{enc(e)}\right)}$. We visualize the raw $768$ dimensional embedding features of prompts, the $10$ dimensional PCA features and the $10$ dimensional deep kernel features for the training split using a two component t-SNE (van der Maaten & Hinton, 2008) in the top row of Figure 3. The $x$- and $y$-axis represent the two t-SNE components, whereas color indicates the validation error of prompts (lighter color indicates better performance). We can see that for both the raw embedding features and the PCA features, it is difficult to visually discern any meaningful clusters or structure how closeness in feature space relates to closeness in performance space. For the deep kernel features, however, we can see that the feature space is well aligned with the performance space (well-performing prompts being closer together with a continuous transition into poorer performing prompts) - of course, this is on the training split on which the GP has been trained on, and therefore these results are not surprising. However, when looking at the test split (that was neither used to perform the PCA nor to train the GP) in the bottom row of Figure 3, we can see that similar conclusions as for the train split hold: The latent representation the feature extractor of the deep kernel has learned during training does generalize to the test split, and it has effectively learned a low-dimensional embedding of prompts aligned with the downstream task.
Figure 3: Visualization of the $768$ dimensional BERT [CLS] token embeddings of prompts via a two component t-SNE. Left: Raw, unprocessed features. Middle: Features of a $10$ component PCA solution. Right: Latent features ($10$ dimensional) from the feature extractor of our structural-aware DK-GP. Top row: Train split. Bottom row: Test split. Color indicates the performance of prompts for LLAMA3 8B Instruct on GSM8K.
@@ Appendix B On the Generalization from Validation to Test
When performing black-box prompt selection, we evaluate prompts on a validation set and iteratively improve over the current best prompt (the incumbent), trying to identify a better one. While progress on the validation set is expected (i.e., if we perform full-fidelity evaluations using the same validation instances for each prompt, the validation error of the incumbent will be monotonically decreasing as optimization progresses), it must not necessarily be the case that we also improve performance on a held-out test set of instances, i.e., the prompt identified as being validation optimal might not necessarily be optimal on the test set.
In this section, we provide additional insights regarding generalization gaps from validation to test performance. Recall that existing methods for black-box prompt selection (e.g., EASE and MIPROv2) are by design not query-efficient but evaluate all prompts on all validation instances or a random subset (e.g., Wu et al. 2024 used sub-sampled validation sets with as few as 20 instances). We now empirically demonstrate that using small random subsets during optimization is not a sensible choice, because this will result in increased variance of the estimate of the validation error, which prevents us from making correct decisions on the validation set. This can result in generalization issues when moving from the validation set to the test set.
We perform the following experiment: We collect the validation and test error (according to the splits described in Table 5) of 250 prompts on the GSM8K task using LLAMA3 8B Instruct as LLM. We vary the number of validation instances used to evaluate the performance of prompts via bootstrapping, using $k=10,50,100,500$ instead of the original $n_{\mathrm{valid}}=1319$ validation instances. Note that to compute the test error, we always use the full test set. In Figure 4, we provide scatter plots of the validation and test errors of the prompts with mean validation errors obtained via bootstrapping using $k=10,50,100,500$ validation instances vs. validation errors obtained on the full validation set of $1319$ instances. We perform $1000$ bootstrap replicates.
@@
We can observe that if we use too few validation instances ($k=10$ but also $k=50$ and $k=100$ to some extent) even if we select the validation optimal prompt, its test error can be far from optimal, because noise dominates the estimate of the validation error.
In Figure 5, we provide box plots of the bootstrapped variance estimates of the mean validation error over prompts when using smaller validation sets.
Figure 5: Box plots of the bootstrapped variance estimates of the mean validation error of 250 prompts evaluated with LLAMA3 8B Instruct on GSM8K varying the number of validation instances used to estimate the mean validation error.
As expected, the variance of the mean validation error can be substantial when using few validation instances. Note that the bootstrap results align with theoretical expectations under the assumption that the point-wise loss (based on exact match) is a binary random variable following a Bernoulli distribution with success probability $p$ (corresponding to a loss of 00). In this case, the average validation error over $n_{\mathrm{valid}}$ instances follows a Binomial distribution, and the variance of the estimated validation error is given by $\nicefrac{{p(1-p)}}{{n_{\mathrm{valid}}}}$. For example, if a prompt has a true “success probability” of $p=0.5$, then using $n_{\mathrm{valid}}=10$ validation instances yields an expected variance of $\nicefrac{{0.5^{2}}}{{10}}=0.025$.
This has serious practical implications, depending on the variation in the performance of prompts on a downstream task. If the true performance of many prompts is similar, we cannot tell them apart based on their estimated validation error, as noise dominates the signal when using too few validation instances. Moreover, for benchmarking methods for black-box prompt selection this is highly relevant as when using too few validation instances, we cannot determine whether a generalization gap from the validation set to the test set results solely from noisy performance estimates or from internal method mechanisms (such as the optimal transport inspired heuristic employed by EASE to only consider examples that are similar to the validation set) which may further result in overfitting to the validation set.
The empirical results we have presented here further provide justification for using a multi-fidelity scheduler over the validation instances during prompt selection, such as Hyperband. Poor-performing prompts can be differentiated using few validation instances, however, well-performing prompts need to be evaluated on larger sets so that one can effectively tell their validation errors apart.